OpenStudy (anonymous):

simplified version of 3 log x – log y + log z

OpenStudy (jdoe0001):

http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif notice the log rules worksheet try using Rule 3 there, for the 1st term

OpenStudy (jdoe0001):

then use Rule 2 there, for the 1st and 2nd terms then use Rule 1 there, to the result

OpenStudy (anonymous):

log cubed over yz

OpenStudy (anonymous):

??

OpenStudy (anonymous):

log x-cubed times z over y

OpenStudy (jdoe0001):

well, close \(\bf 3 log (x) - log (y) + log(z) \implies log(x^3)- log (y) + log(z)\\ \quad \\ log\left(\cfrac{x^3}{y}\right) + log(z)\implies log\left(\cfrac{x^3}{y}\times z\right) \implies log\left(\cfrac{x^3z}{y}\right) \)

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

Thank you

OpenStudy (jdoe0001):

yw